OpenStudy (anonymous):

How do you solve this radical equation: (x+6)^1/2 +6=x?

OpenStudy (anonymous):

\[\sqrt{x+6}+6=x\] \[\sqrt{x+6}=x-6\] squaring on both sides \[(\sqrt{x+6})^2=(x-6)^2\] \[x+6=x^2-12x+36\] \[0=x^2-12x-x+36-6\] \[x^2-13x+30=0\] \[x^2-10x-3x+30=0\] \[x(x-10)-3(x-10)\] \[(x-3)(x-10)=0\] \[x=3,x=10\]